# MDPs with a State Sensing Cost

### v1

@@ 6 Concluding remarks
In this paper, we have analysed a class of MDPs with a state sensing cost. Here, the agent must, in a history-dependent, opportunistic manner, determine when to sense the state of the system/environment. While these MDPs are intractable under generic planning algorithms, we exploit the special structure of these sensing-cost MDPs to devise intelligent heuristics and truncation approaches with provable optimality/suboptimality guarantees.
@@
At a high level, this work is related to the vast recent literature on Age of Information (AoI), where the goal is to allocate resources or incur costs so as to minimize the age (a.k.a., staleness) of the state information; see (Yates et al., 2021) for a survey. However, the AoI literature does not, as such, consider the control aspect where the goal of state estimation is actually to influence the state evolution favourably. Additionally, the AoI literature employs a universal (state-independent) age/staleness penalty; in practice, one would expect that the agent would be more tolerant of delayed state information in certain states than others. The present formulation seeks to formally capture this trade-off.
@@ Appendix A Proofs
Let $\pi$ be the AS policy for the MDP $\mathcal{M}_{k}$, then for all states $\tilde{s}\in\mathcal{S}_{\infty}$, $\pi(\tilde{s})=(\pi_{AS}(\tilde{s}),sense)$. We now apply policy improvement on $\pi$. Note that policy $\pi$ cannot be improved by any sensing action for any of the states, i.e.,
$Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a,sense))\geq V^{\pi}_{\mathcal{M}_{k}}(\tilde{s})\quad\forall\tilde{s}\in\mathcal{S}_{\infty},\ a\in A$
. Furthermore, if (4) holds, then policy $\pi$ cannot be improved by taking blind actions and is therefore optimal.
$Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a_{1},blind))$ $\geq Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a_{1},sense)),$ (4)
$\quad\forall(\tilde{s},a_{1})\in\mathcal{S}_{\infty}\times A.$
@@
$\mathcal{B}(\tilde{s})\mathcal{C}(a_{1})+\alpha\left(k+\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2})\right)\\ +\alpha^{2}\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\pi}_{\mathcal{M}_{k}}\\ \geq\mathcal{B}(\tilde{s})\mathcal{C}(a_{1})+k+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\pi}_{\mathcal{M}_{k}},$
@@
$\frac{k}{1-\alpha}+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\ast}\\ \geq\frac{k}{\alpha(1-\alpha)}+\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\ast}-\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2}).$
@@
$\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2})-\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\ast}\\ +\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\ast}\geq\frac{k}{\alpha}.$
@@
$V_{\mathcal{M}_{k,N}}^{\pi}(i)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}V_{\mathcal{M}_{k,N}}^{\pi}(j)\Biggr{)},$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(i)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}V_{\mathcal{M}_{k,N}}^{*}(j)\Biggr{)}.$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)=$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
$\implies$ $\,V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)\leq$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
[…]
@@
$B_{\pi}^{N}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq$
$\quad\alpha^{N}(V_{\mathcal{M}_{k,N}}^{*}(s_{N})-V_{\mathcal{M}_{k}}^{\pi}(s_{N}))\leq\frac{\alpha^{N}k}{1-\alpha},$
@@
where $s_{N}\in\mathcal{L}^{j}_{N}$, is the state reached after playing $N$ blind steps starting from $j$ following $\pi$. This establishes (8). ∎
@@
Define $\pi_{N+1}$ as an extension of the policy $\pi^{*}_{\mathcal{M}_{k,N}}$ for $\mathcal{M}_{k,N+1}$. Without loss of generality (W.L.O.G.), assume that $\pi^{*}_{\mathcal{M}_{k,N}}(\tilde{s})\in A_{s}$ for all states $\tilde{s}\in\mathcal{L}_{N}$. Under $\pi_{N+1}$, states $\tilde{s}\in\bigcup_{l=0,1,\ldots,N}\mathcal{L}_{l}$ (i.e., all states from layers 00 to $N$) are mapped to actions provided by the policy $\pi^{*}_{\mathcal{M}_{k,N}}$, while states $\tilde{s}\in\mathcal{L}_{N+1}$ (i.e., states in the $N+1$ layer) are assigned arbitrary actions.
@@
for $10\leq m\leq T-$ 1, where the only difference from $Z(s_{T})$ is that we start from state $s_{m}$ at $t=0$ and calculate the cumulative cost to reach $s_{T}$.
$Z(i)+\alpha^{N+1}\min_{a}\Bigl{(}\mathcal{B}(i)\mathcal{C}(a)+k+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}}\Bigr{)}\\ \quad\quad\geq V^{*}_{\mathcal{M}_{k,N}}(j)\quad\forall j\in\mathcal{S},\ i\in\mathcal{L}^{j}_{N+1}.$ (10)
@@
$Z_{\tilde{s}}(i)+\alpha^{N+1-m}\left(\mathcal{B}(i)\mathcal{C}(a)+k+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}}\right)\\ <V^{\pi^{\prime}_{N+1}}_{\mathcal{M}_{k,N}}(\tilde{s})$ (11)
@@
$Z(\tilde{s})+Z_{\tilde{s}}(i)+\alpha^{N+1-m}(\mathcal{B}(i)\mathcal{C}(a)+k\\ \quad+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}})\geq Z(\tilde{s})+V^{*}_{\mathcal{M}_{k,N}}(\tilde{s}).$
@@
$Z_{\tilde{s}}(i)+\alpha^{N+1-m}(\mathcal{B}(i)\mathcal{C}(a)+k\\ +\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}})\geq V^{*}_{\mathcal{M}_{k,N}}(\tilde{s}),$
@@
$V^{*}_{\mathcal{M}_{k,N}}(j)\leq\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right)\\ \leq V_{\mathcal{M}_{0}}^{\pi_{j}^{M}}(j)\leq V_{\mathcal{M}_{k}}^{\pi_{j}^{M}}(j)\\ \implies V^{*}_{\mathcal{M}_{k}}(j)\leq V^{\pi^{*}_{\mathcal{M}_{k,N}}}_{\mathcal{M}_{k}}(j)=V^{*}_{\mathcal{M}_{k,N}}(j)\leq V_{\mathcal{M}_{k}}^{\pi_{j}^{M}}(j)$
@@
$V_{\mathcal{M}_{k,N}}^{\pi}(l)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}V_{\mathcal{M}_{k,N}}^{\pi}(j)\Biggr{)},$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(l)=$ $Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))$
$\quad+\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}V_{\mathcal{M}_{k,N}}^{*}(j)\Biggr{)}.$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(l)-V_{\mathcal{M}_{k}}^{\pi}(l)=$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
$\implies$ $\,V_{\mathcal{M}_{k,N}}^{*}(l)-V_{\mathcal{M}_{k}}^{\pi}(l)\leq$
$\quad\alpha^{m+1}\Biggl{(}\sum_{j\in\mathcal{L}_{0}}p_{lj}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Biggr{)}$
@@
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ $\leq\alpha^{a_{i}}\Bigl{(}\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl{(}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr{)}\Bigr{)},$
$\quad\forall i\in\bar{G},$
@@ 2. It follows from the same proof that the stronger claim below holds for any root state $j$:
$V^{*}_{\mathcal{M}_{k}}(j)\geq\min\Bigl{\{}\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right),V^{*}_{\mathcal{M}_{k,N}}(j)\Bigr{.}\\ \Bigl{.}-\alpha\max_{s\in\mathcal{S}\setminus\{j\}}\left[V^{*}_{\mathcal{M}_{k,N}}(s)-\min_{r\in\mathcal{L}^{s}_{N+1}}\left(Z(r)+\alpha^{N+1}V_{AS;0}(r)\right)\right]^{+}\Bigr{\}}.$
Here, $[x]^{+}$ denotes the positive part of $x.$
@@
$V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq V_{\mathcal{M}_{k,N}}^{*}(j)\\ -\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right).$
A.5 Proof of Lemma 4.5
@@
$V_{AS;0}(\tilde{s})$ $=\mathcal{B}(\tilde{s})\mathcal{C}(\pi_{AS}(\tilde{s}))+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(\pi_{AS}(\tilde{s}))V^{*},$ (14)
where $\mathcal{B}(\tilde{s})=\mathcal{B}(\tilde{s})\mathcal{T}(\pi_{AS}(\tilde{s}))$. We claim that for the above non-trivial MDP, for each $a\in A$, there exists a root state $r$ such that
@@
Also note that $\exists\ N^{*}$ s.t. $\forall N\geq|\mathcal{S}|-1$, every element of $\mathcal{B}(\tilde{s})$ is non-zero $\forall\tilde{s}\in\mathcal{L}_{N}$. Hence, by applying the inequality from (15), we obtain
@@
$\implies\mathcal{B}(\tilde{s})V^{*}$ $<V_{AS;0}(\tilde{s}).$ (16)
Thus, for all $\tilde{s}\in\mathcal{L}_{N}$, where $N\geq|\mathcal{S}|-2$, applying (16) to the definition of $V_{AS;0}(\tilde{s})$ in (14), we obtain
@@
$Z(i)+\alpha^{N}V_{AS;0}(i)<Z(i)+\alpha^{N}\left(\mathcal{B}(i)\mathcal{C}(a)+\alpha V_{AS;0}(i^{\prime})\right)\\ =Z(i^{\prime})+\alpha^{N+1}V_{AS;0}(i^{\prime}).$

### v3

@@ 6 Discussion and Conclusion
In this paper, we have analysed a class of MDPs with a state sensing cost. Here, the agent must, in a history-dependent, opportunistic manner, determine when to sense the state of the system. While these MDPs are intractable under generic planning algorithms, we exploit the special structure of these sensing-cost MDPs to devise clever algorithms and truncation approaches with provable optimality/suboptimality guarantees. Our main contributions are summarized as follows:
$\bullet$ We reformulate opportunistic state sensing in an expanded state-space framework and propose a novel planning algorithm, SPI. SPI outperforms state-of-the-art POMDP solvers and specialized solvers for this setting (e.g., $\pi_{\text{ATM}}$) across diverse domains for all ranges of sensing costs within reasonable compute time (Table 1). Moreover, it scales to large-state-space MDPs and naturally extends to continuous state spaces using function approximation (Appendix C.2).
$\bullet$ Using truncated MDP analysis, we derive lower bounds on the optimal value function, which enable explicit computation of suboptimality gaps for any policy (Theorems 4.4 and 4.2).We establish a quantitative threshold on the sensing cost (Theorem 4.1) below which always sensing is provably optimal.
To the best of our knowledge, this work presents the first set of results establishing suboptimality guarantees for policies and explicit cost thresholds in this setting.
In practice, sensing costs are often context-dependent, shaped by factors such as resource constraints or belief uncertainty (e.g., in medical decision-making or robotics). Our framework is readily extendable to non-uniform sensing costs: both SPI and the truncated MDP approach, along with the theoretical results, apply beyond the constant-$k$ setting (see Appendix F).
@@
At a high level, this work is related to the vast recent literature on Age of Information (AoI), where the goal is to allocate resources or incur costs so as to minimize the age (a.k.a., staleness) of the state information; see Yates et al. (2021) for a survey. However, the AoI literature does not, as such, consider the control aspect where the goal of state estimation is actually to influence the state evolution favourably. Additionally, the AoI literature employs a universal (state-independent) age/staleness penalty; in practice, one would expect that the agent would be more tolerant of delayed state information in certain states than others. The present formulation seeks to formally capture this trade-off.
Appendix
Appendix Contents
A. Proofs ........................................................................................................................................................................A
A.1 Proof of Theorem 4.1 ........................................................................................................................................................................A.1
A.2 Proof of Theorem 4.2 ........................................................................................................................................................................A.2
A.3 Proof of Lemma 4.3 ........................................................................................................................................................................A.3
A.4 Proof of Theorem 4.4 ........................................................................................................................................................................A.4
A.5 Proof of Lemma E.1 ........................................................................................................................................................................A.5
B. Experimental Design and Setup ........................................................................................................................................................................B
C. Analysis of SPI ........................................................................................................................................................................C
D. Analysis of ATM Heuristic ........................................................................................................................................................................D
E. Additional Results ........................................................................................................................................................................E
F. Extension to Non-Uniform Sensing Costs ........................................................................................................................................................................F
G. Inventory Management Case Study ........................................................................................................................................................................G
@@ Appendix A Proofs
We define $\pi$ to be the AS policy for the MDP $\mathcal{M}_{k}$ and hence for each $\tilde{s}\in\mathcal{S}_{\infty},\pi(\tilde{s})=(\pi_{AS}(\tilde{s}),sense)$. It is important to note that AS policy $\pi_{AS}$ takes the optimal sensing action in each belief state and agrees with the optimal policy $\pi^{*}$ associated with the baseline MDP $\mathcal{M}$ over root states. Therefore, we have:
[…]
$=\min_{a\in A}\left(B(\tilde{s})\mathcal{C}(a)+\alpha B(\tilde{s})\mathcal{T}(a)V_{\mathcal{M}_{k}}^{\pi}\right)+k\quad(\text{Since }V_{\mathcal{M}_{k}}^{\pi}(s)=V^{*}(s)+\frac{k}{1-\alpha}\quad\forall s\in\mathcal{S})$
$=\min_{a\in A}Q_{\mathcal{M}_{k}}^{\pi}(\tilde{s},(a,sense))$
Hence, policy $\pi$ cannot be improved by any sensing action for any of the states, i.e.,
$Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a,sense))\geq V^{\pi}_{\mathcal{M}_{k}}(\tilde{s})\quad\forall\tilde{s}\in\mathcal{S}_{\infty},\ a\in A.$
$Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a_{1},blind))$ $\geq Q^{\pi}_{\mathcal{M}_{k}}(\tilde{s},(a_{1},sense))\quad\forall(\tilde{s},a_{1})\in\mathcal{S}_{\infty}\times A.$ (3)
[…]
@@
$\mathcal{B}(\tilde{s})\mathcal{C}(a_{1})+\alpha\left(k+\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2})\right)+\alpha^{2}\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\pi}_{\mathcal{M}_{k}}\geq\mathcal{B}(\tilde{s})\mathcal{C}(a_{1})+k+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\pi}_{\mathcal{M}_{k}},$
@@
$\frac{k}{1-\alpha}+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\ast}\geq\frac{k}{\alpha(1-\alpha)}+\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\ast}-\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2}).$
@@
$\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{C}(a_{2})-\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})V^{\ast}+\alpha\mathcal{B}(\tilde{s})\mathcal{T}(a_{1})\mathcal{T}(a_{2})V^{\ast}\geq\frac{k}{\alpha}.$
@@
$V_{\mathcal{M}_{k,N}}^{\pi}(i)=\$$Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))+\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{ij}V_{\mathcal{M}_{k,N}}^{\pi}(j)\Biggr),$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(i)=\$$Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))+\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{ij}V_{\mathcal{M}_{k,N}}^{*}(j)\Biggr).$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)=\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl(V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr)\Biggr)$
$\implies$ $\,V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)\leq\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl(V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr)\Biggr)$
$\text{(since $V_{\mathcal{M}_{k,N}}^{*}\leq B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}$ elementwise)}.$
@@
$B_{\pi}^{N}V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq\alpha^{N}(V_{\mathcal{M}_{k,N}}^{*}(s_{N})-V_{\mathcal{M}_{k}}^{\pi}(s_{N}))\leq\frac{\alpha^{N}k}{1-\alpha},$
@@
where $s_{N}\in\mathcal{L}^{j}_{N}$, is the state reached after playing NN blind steps starting from $j$ following $\pi$. This establishes (7). ∎
@@
Define $\pi_{N+1}$ as an extension of the policy $\pi^{*}_{\mathcal{M}_{k,N}}$ for $\mathcal{M}_{k,N+1}$. Without loss of generality (W.L.O.G.), assume that $\pi^{*}_{\mathcal{M}_{k,N}}(\tilde{s})\in A_{s}$ for all states $\tilde{s}\in\mathcal{L}_{N}$. Under $\pi_{N+1}$, states $\tilde{s}\in\bigcup_{l=0,1,\ldots,N}\mathcal{L}_{l}$ (i.e., all states from layers 0 to $N$) are mapped to actions provided by the policy $\pi^{*}_{\mathcal{M}_{k,N}}$, while states $\tilde{s}\in\mathcal{L}_{N+1}$ (i.e., states in the $N+1$ layer) are assigned arbitrary actions.
@@
for $0\leq m\leq T-1$, where the only difference from $Z(s_{T})$ is that we start from state $s_{m}$ at $t=0$ and calculate the cumulative cost to reach $s_{T}$.
$Z(i)+\alpha^{N+1}\min_{a}\Bigl(\mathcal{B}(i)\mathcal{C}(a)+k+\alpha\mathcal{B}(i)\mathcal{T}(a)V^{*}_{\mathcal{M}_{k,N}}\Bigr)\geq V^{*}_{\mathcal{M}_{k,N}}(j)\quad\forall j\in\mathcal{S},\ i\in\mathcal{L}^{j}_{N+1}.$ (9)
@@
[…]
@@
[…]
@@
[…]
@@
$V^{*}_{\mathcal{M}_{k,N}}(j)$ $\leq\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right)\leq V_{\mathcal{M}_{0}}^{\pi_{j}^{M}}(j)\leq V_{\mathcal{M}_{k}}^{\pi_{j}^{M}}(j)$
$\implies V^{*}_{\mathcal{M}_{k}}(j)$ $\leq V^{\pi^{*}_{\mathcal{M}_{k,N}}}_{\mathcal{M}_{k}}(j)=V^{*}_{\mathcal{M}_{k,N}}(j)\leq V_{\mathcal{M}_{k}}^{\pi_{j}^{M}}(j)$
@@
$V_{\mathcal{M}_{k,N}}^{\pi}(l)=\$$Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))+\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{lj}V_{\mathcal{M}_{k,N}}^{\pi}(j)\Biggr),$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(l)=\$$Z(s_{m})+\alpha^{m}\mathcal{C}_{N}(s_{m},\pi(s_{m}))+\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{lj}V_{\mathcal{M}_{k,N}}^{*}(j)\Biggr).$
@@
$B_{\pi}^{m+1}V_{\mathcal{M}_{k,N}}^{*}(l)-V_{\mathcal{M}_{k}}^{\pi}(l)=\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{lj}\bigl(V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr)\Biggr)$
$\implies$ $\,V_{\mathcal{M}_{k,N}}^{*}(l)-V_{\mathcal{M}_{k}}^{\pi}(l)\leq\alpha^{m+1}\Biggl(\sum_{j\in\mathcal{L}_{0}}p_{lj}\bigl(V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr)\Biggr)$
@@
$V_{\mathcal{M}_{k,N}}^{*}(i)-V_{\mathcal{M}_{k}}^{\pi}(i)$ $\leq\alpha^{a_{i}}\Bigl(\sum_{j\in\mathcal{L}_{0}}p_{ij}\bigl(V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\bigr)\Bigr),\quad\forall i\in\bar{G},$
@@ 2. It follows from the same proof that the stronger claim below holds for any root state $j$:
$\hskip-8.0ptV^{*}_{\mathcal{M}_{k}}(j)\geq\min\Bigl\{\min_{i\in\mathcal{L}^{j}_{N+1}}\hskip-6.0pt\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right),V^{*}_{\mathcal{M}_{k,N}}(j)-\alpha\hskip-2.0pt\max_{s\in\mathcal{S}\setminus\{j\}}\left[V^{*}_{\mathcal{M}_{k,N}}(s)-\hskip-6.0pt\min_{r\in\mathcal{L}^{s}_{N+1}}\hskip-5.0pt\left(Z(r)+\alpha^{N+1}V_{AS;0}(r)\right)\right]^{+}\Bigr\}.$
Here, $x]^{+}$ denotes the positive part of $x.$
@@
$V_{\mathcal{M}_{k,N}}^{*}(j)-V_{\mathcal{M}_{k}}^{\pi}(j)\leq V_{\mathcal{M}_{k,N}}^{*}(j)-\min_{i\in\mathcal{L}^{j}_{N+1}}\left(Z(i)+\alpha^{N+1}V_{AS;0}(i)\right).$
A.5 Proof of Lemma E.1
@@
[…]
where $)\mathcal{B}(\tilde{s})=\mathcal{B}(\tilde{s})\mathcal{T}(\pi_{AS}(\tilde{s}$)). We claim that for the above non-trivial MDP, for each $a\in A$, there exists a root state $r$ such that
@@
[…]
@@
[…]
@@
$Z(i)+\alpha^{N}V_{AS;0}(i)<Z(i)+\alpha^{N}\left(\mathcal{B}(i)\mathcal{C}(a)+\alpha V_{AS;0}(i^{\prime})\right)=Z(i^{\prime})+\alpha^{N+1}V_{AS;0}(i^{\prime}).$